OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) >= -31. Can someone please tell me how to solve the first part and that's all. Do I combine like terms or what??

jimthompson5910 (jim_thompson5910):

8x - (5x + 4) >= -31 ... Start with the given inequality. 8x - 5x - 4 >= -31 ... Distribute 3x - 4 >= -31 ... Combine like terms. 3x - 4 +4 >= -31 +4 ... Add 4 to both sides. 3x >= -27 ... Combine like terms. 3x/3 >= -27/3 ... Divide both sides by 3 to completely isolate x. x >= -9 ... Reduce